Rhino viruses typically cause common colds. In a test of the effectiveness of echinacea,

3838

of the

4343

subjects treated with echinacea developed rhinovirus infections. In a placebo group,

8484

of the

9999

subjects developed rhinovirus infections. Use a

0.010.01

significance level to test the claim that echinacea has an effect on rhinovirus infections. Complete parts (a) through (c) below.

z value

pvalue
99% confidence interval
Answer #1

Answer:

Rhino viruses typically cause common colds. In a test of the effectiveness of echinacea,

38 of the43 subjects treated with echinacea developed rhinovirus infections. In a placebo group,

84 of the 99 subjects developed rhinovirus infections. Use a0.01significance level to test the claim that echinacea has an effect on rhinovirus infections. Complete parts (a) through (c) below.

H_0: P_1 = P_2 \ \ \ H_1: P_1 > P_2

Upper tail test used.

z value=0.5546

p value=0.2896

99% confidence interval (-0.1212, 0.1917)

calculations:

z = \frac{ p_1 - p_2}{\sqrt{ P*(1-P)(1/n_1 + 1/n_2) }}

Z Test for Differences in Two Proportions

Data

Hypothesized Difference

0

Level of Significance

0.01

Group 1

Number of Items of Interest

38

Sample Size

43

Group 2

Number of Items of Interest

84

Sample Size

99

Intermediate Calculations

Group 1 Proportion

0.8837

Group 2 Proportion

0.8485

Difference in Two Proportions

0.0352

Average Proportion

0.8592

Z Test Statistic

0.5546

Upper-Tail Test

Upper Critical Value

2.3263

p-Value

0.2896

Do not reject the null hypothesis

CI = p1-p2 ± z*se

Standard error of p1-p2= se=

/P1 * (1-p1)/m +Pa * (1-p2) /n2

Confidence Interval Estimate

of the Difference Between Two Proportions

Data

Confidence Level

99%

Intermediate Calculations

Z Value

2.576

Std. Error of the Diff. between two Proportions

0.0607

Interval Half Width

0.1564

Confidence Interval

Interval Lower Limit

-0.1212

Interval Upper Limit

0.1917